OpenStudy (anonymous):

Hi! Could someone pleaseeeeeeee help me!! If the legs of a right triangle have lengths of 6 inches and 8 inches, what is the length of the hypotenuse in inches? 11 13 10 16 I have 3 other questions like this, and i need someone to explain how to do this so i can answer them all correctly while understanding!

OpenStudy (anonymous):

Pythagorean theorem is: A^2 + B^2 = C^2 so: 6^2 + 8^2 then take the sqrt of that and there is your answer

OpenStudy (anonymous):



OpenStudy (anonymous):

Use Pythagoras theorem. Take AB = 8 inches BC = 6 inches AC = ? So, According to pythagoras theorem, AB^2 + BC^2 = AC^2 So, put their values, 8^2 + 6^2 = AC^2 64 + 36 = 100 = AC^2 Take square root of 100, which is 10. So, AC = 10 inches So your answer is 10.

hero (hero):

What happens if you multiply every side of a 3:4:5 triangle by 2?